OpenStudy (anonymous):

evaluate the arc length of the curve corresponding to the following parametric equations : x(t) = 4 sin (4t) y(t)=4cos (4t) and it's the full circle which is from 0 to 2pi my solution : 32pi , is it correct ?

OpenStudy (anonymous):

depends

OpenStudy (anonymous):

from \(0\) to \(2\pi\) is more that a circle it goes around the circle 4 times

OpenStudy (anonymous):

i guess on rotation will give you \(8\pi\) so four will give you \(32\pi\) looks good in that case

OpenStudy (anonymous):

Thanks man.